# Self-Supervised Learning for Time Series Analysis: Taxonomy, Progress, and Prospects

### v1

@@ 3.3 Diffusion-based generation
As a new kind of deep generative model, diffusion models have achieved great success recently in many fields, including image synthesis, video generation, speech generation, bioinformatics, and natural language processing due to their powerful generating ability [85, 86, 87, 88, 89]. We provide a brief introduction to this class of methods and show the main applications in time series analysis recently.
The key design of the diffusion model contains two inverse processes: the forward process of injecting random noise to destruct data and the reverse process of sample generation from noise distribution (usually normal distribution). The intuition is that if the forward process is done step-by-step with a transition kernel between any two adjacent states, then the reverse process can follow a reverse state transition operation to generate samples from noise (the final state of the forward process). However, it is usually not easy to formulate the reverse transition kernel, and thus diffusion models learn to approximate the kernel by deep neural networks. Nowadays, there are mainly three basic formulations of diffusion models: denoising diffusion probabilistic models (DDPMs) [85, 90], score matching diffusion models [91, 92], and score SDEs [93, 94].
3.3.1 DDPMs
The forward and reverse processes in DDPMs are two Markov chains: a forward chain that adds random noise to data and a reverse chain that transforms noise back into data. Formally, denote the data distribution as $\boldsymbol{x}_{0}\sim q(\boldsymbol{x}_{0})$, the forward Markov process gradually adds Gaussian noise to the data according to transition kernel $q(\boldsymbol{x}_{t}|\boldsymbol{x}_{t-1})$. It generates a sequence of random variables $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots,\boldsymbol{x}_{T}$. Thus the joint distribution of $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots,\boldsymbol{x}_{T}$ conditioned on $\boldsymbol{x}_{0}$ is
@@ 3.3.1 DDPMs
where $\theta$ is the model parameters and $p_{\theta}(\boldsymbol{x_{t-1}|x_{t}})=\mathcal{N}(\boldsymbol{x}_{t-1};\mu_{\theta}(\boldsymbol{x}_{t},t),\sum_{\theta}(\boldsymbol{x}_{t},t)).$ The key to achieving the success of sample generating is training the parameters $\theta$ to match the actual reverse process, that is, minimizing the Kullback-Leibler divergence between the two joint distributions. Thus, the training loss is
@@
$\geq\mathbf{E}[-\log p_{\theta}(\boldsymbol{x}_{0})]+const,$
according to Jensen’s inequality. See more details in [87].
3.3.2 Score matching diffusion models
For a given probability density function $p(\boldsymbol{x})$, the Stein score (noted as $\nabla_{x}\log p(\boldsymbol{x})$), which is the gradient of the log-density function of the input data $\boldsymbol{x}$, points to the directions with the largest growth rate in probability density function. The key idea of score based diffusion models is to perturb data with a sequence of Gaussian noise and then jointly estimate the score functions for all noisy data distributions by training a deep neural network conditioned on noise levels. The motivation of the idea is that, in many situations, it is easier to model and estimate the score function than the original probability density function.
Langevin dynamics is one of the proper techniques. With a step size $\alpha>0$, a number of iterations $T$ and an initial sample $x_{0}$, Langevin dynamics iteratively does the following estimation to gain a close approximation of $p(\boldsymbol{x})$
@@ 3.3.2 Score matching diffusion models
3.3.3 Score SDEs
DDPMs and SGMs are both limited to discrete and finite time steps (which can also be called noise levels). Score SDEs process the diffusion operation according to the stochastic differential equation (SDE) [94]:
@@ 3.3.3 Score SDEs
where $\delta(\frac{t}{T})=\delta_{t}$ as T goes to infinity.
With any diffusion process in the form of (16), the reverse process can be gained by solving the following SDE:
@@
Diffusion models have also been applied in time series analysis recently. We will briefly introduce some of them, including the designed architectures and the main diffusion techniques used.
3.3.4 Diffusion models for time series imputation
Conditional Score-based Diffusion Models for Imputation (CSDI) [95] proposes a novel time series imputation method that utilizes score-based diffusion models conditioned on observed data. It is trained for imputation and can be easily generalized to forecasting by imputing data at the end of the time series sequence. To deal with the issue that the ground-truth missing values are not accessible in practice, a self-supervised training procedure is also proposed where values are masked as missing values to train the whole architecture.
3.3.5 Diffusion models for time series forecasting
Several works have appeared that take advantage of diffusion models to forecast time series. TimeGrad [96] takes an RNN conditioned diffusion probabilistic model at some time step to depict the fixed forward process and the learned reverse process. The cross-correlation of the time series is also depicted during the process stages. It is an autoregressive model for multivariate time series forecasting and learns gradients by optimizing the parameters that minimize the negative log-likelihood of the model:
$\sum_{t=t_{0}}^{T}-\log p_{\theta}(x_{t}^{0}|\boldsymbol{h}_{t-1}),$ (21)
@@ 3.3.5 Diffusion models for time series forecasting
4 Contrastive-based Methods
Contrastive learning is a widely used self-supervised learning strategy, showing a strong learning ability in computer vision and natural language processing. Unlike discriminative models that learn a mapping rule to true labels and generative models that try to reconstruct inputs, contrastive-based methods aim to learn data representations by contrasting between positive and negative samples. Specifically, positive samples should have similar representations, while negative samples have different representations. Therefore, the selection of positive samples and negative samples is very important to contrastive-based methods. This section sorts out and summarizes the existing contrastive-based methods in time series modeling according to the selection of positive and negative samples.
@@ 4.1 Sampling contrast
where $S(\cdot)$ denotes the similarity of the two representations. However, due to the non-stationary characteristics of most time series data, it is still a challenge to choose the correct positive and negative samples based on contextual information in time series data. Temporal neighborhood coding (TNC) was recently proposed to deal with this problem [102]. TNC uses Augmented Dickey-Fuller (ADF) statistical test to determine the stationary region and introduces PU (Positive-Unlabeled) learning to handle the problem of sampling bias by treating negative samples as unknown samples and then assigning weights to these samples. The learning objective is extended to
@@
where $w$ is the probability of sampling false negative samples. $\mathcal{N}$ denotes the neighboring area and $\tilde{\mathcal{N}}$ denotes the non-neighboring area. Supervised contrastive learning (SCL) [103] effectively addresses sampling bias, so introducing the supervised signal to identify positive and negative samples is a feasible solution. Neighborhood contrastive learning (NCL) is a recent time series modeling method that combines context sampling and the supervised signal to generate positive and negative samples [104]. NCL assumes that if two samples share some predefined attributes, then they are considered to share the same neighboring area.
@@ 4.2 Prediction contrast
In this category, prediction tasks that use the context (present) to predict the target (future information) are considered self-supervised pretext tasks, and the goal is to maximally preserve the mutual information of the context and the target. Contrastive predictive coding (CPC) proposed by [22] provides a contrastive learning framework to perform the prediction task using InfoNCE loss. As shown in Figure 4(b), the context $c_{t}$ and the sample from $p(x_{t+k}|c_{t})$ constructs positive pairs, and the samples form the ‘proposal’ distribution $p(x_{t+k})$ are negative samples. The learning objective is as follows:
@@ 4.3 Augmentation contrast
Since time series data can be converted to frequency domain representations through Fourier transform, the augmentation method can be developed from the time and frequency domains. In the time domain, TS-TCC[111] and its extended version CA-TCC[112] designed two time series data augmentation techniques, one is strong augmentation (Permutation-and-Jitter), and the other is weak augmentation (Jitter-and-Scale). TS2Vec [113] generates different views through masking operations that randomly mask out some time steps. Generally speaking, there is no one-size-fits-all answer to the choice of data augmentation methods. Therefore, some works comprehensively compare and study the augmentation methods and further evaluate the performance on different tasks [114, 33, 115, 116].
Data augmentation in the frequency domain is also feasible for time series data. CoST [117] is a disentangled seasonal-trend representation learning method, which uses Fast Fourier Transform to convert different augmented views into amplitude and phase representations, and then uses (26) to train the model. BTSF [118] is a contrastive-based method based on a time-frequency fusion strategy, which first generates an augmented view in the time domain through the dropout operation and then generates another augmented view in the frequency domain through Fourier transform. Finally, the bilinear temporal-spectral fusion mechanism is used to achieve the fusion of time-frequency information. However, CoST and BTSF do not directly modify the frequency representation, while TF-C [119] is the first work to augment the time series data through frequency perturbations, which has achieved better performance than TS2Vec [113] and TS-TCC [111]. TF-C implements three augmentation strategies: Low- vs. High-band perturbations, Single- vs. Multi-component perturbations, and Random vs. Distributional perturbations.
@@ 4.4 Prototype contrast
In time series modeling based on prototypes contrast, ShapeNet [130] takes Shaplets as input and constructs a cluster-level triplet loss, which considers the distance between the anchor and multiple positive (negative) samples as well as the distance between positive (negative) samples. ShapeNet is an implicit prototype contrast because it does not introduce explicit prototypes (cluster centers) during the training phase. TapNet [131] and VSML [132] are explicit prototypes contrast because explicit prototypes are introduced. TapNet introduces a learnable prototype for each predefined class and classifies the input time series sample according to the distance between the sample and each class prototype. VSML defines virtual sequences, which have the same function as prototypes, i.e., minimize the distance between samples and virtual sequences, but maximize the distance between virtual sequences.
MHCCL [133] proposes a hierarchical clustering based on the upward masking strategy and a contrastive pairs selection strategy based on the downward masking strategy. In the upward mask strategy, MHCCL believes that outliers greatly impact prototypes, so these outliers should be removed when updating prototypes. The downward masking strategy, in turn, uses the clustering results to select positive and negative samples, i.e., samples belonging to the same prototype are regarded as true positive samples, and samples belonging to different prototypes are regarded as true negative samples.
@@ 4.5 Expert knowledge contrast
Expert knowledge contrast is a relatively new representation learning framework. Generally speaking, this modeling framework incorporates expert prior knowledge or information into deep neural networks to guide model training [134, 135]. In the contrastive learning framework, prior knowledge can help the model choose the correct positive and negative samples during training. An example of expert knowledge contrast is shown in Fig. 4(e). Here we sort out three interesting works.
Shi et al. [136] used the DTW distance of time series samples as prior information and believed that two samples with a small DTW distance have a higher similarity. Specifically, given the anchor $x^{ref}$ and the other two samples $x_{i}$ and $x_{j}$, the DTW distance between $x^{ref}$ and the other two samples is calculated first, then the sample with a small distance from $x^{ref}$ is considered as the positive sample of $x^{ref}$. This selection process is defined as
@@ 5 Adversarial-based Methods
According to the final task, the existing adversarial-based representation learning methods can be divided into time series generation and imputation, and auxiliary representation enhancement. The examples of these adversarial-based SSL are shown in Fig. 5.
@@ 6 Applications and Datasets
TABLE II: Summary of time series applications and widely used datasets. The UCR and UEA datasets have multiple sub-datasets respectively, so their size and dimension are not fixed values. The size and dimension of each sub-dataset are represented by $\mathcal{M}$ and $\mathcal{D}$, respectively. AnRa represents Anomaly Ratio. SaIn represents Sampling Interval.
Applications Dataset Size Dimension Source Citations Comment
@@
SSL has many applications across different time series tasks, such as forecasting, classification, anomaly detection, etc. We summarize the most widely used datasets and representative references according to the application area. In this article, we collect widely used datasets in four tasks: anomaly detection, forecasting, classification, and clustering. As shown in Table II, we provide useful information, including dataset name, dimension, size, source, citations, and useful comments. Datasets for classification tasks and clustering tasks are listed together because the goals of these two tasks are consistent.
@@
Anomaly Detection. The anomaly detection problem for time series is usually formulated as identifying outlier time points or unexpected time sequences relative to some norm or usual signal. Most time series anomaly detection methods are constructed under an unsupervised learning framework because obtaining labels for anomalous data is challenging. Autoregressive-based forecasting and autoencoder-based reconstruction are the most commonly used modeling strategies. To be concrete, THOC [51] and GDN [54] employ autoregressive-based forecasting SSL framework, which assumes that anomalous sequences or time points are not predictable. While based on the principle that anomalous are hard to reconstruct, RANSynCoders [60], USAD [63], AnomalyTrans [154], and DAEMON [165] employ autoencoder-based reconstruction SSL framework. Furthermore, VGCRN [80] and FuSAGNet [64] combine two frameworks to achieve more robust and accurate detection results. It is beneficial to introduce an adversarial-based SSL method in the detection models, which can further amplify the difference between normal and anomalous data, such as USAD [63] and DAEMON [165].

### v3

@@ 3.3 Diffusion-based generation
As a new kind of deep generative model, diffusion models have achieved great success recently in many fields, including image synthesis, video generation, speech generation, bioinformatics, and natural language processing due to their powerful generating ability [88, 89, 90, 91, 92]. The key design of the diffusion model contains two inverse processes: the forward process of injecting random noise to destruct data and the reverse process of sample generation from noise distribution (usually normal distribution). The intuition is that if the forward process is done step-by-step with a transition kernel between any two adjacent states, then the reverse process can follow a reverse state transition operation to generate samples from noise (the final state of the forward process). However, it is usually not easy to formulate the reverse transition kernel, and thus diffusion models learn to approximate the kernel by deep neural networks. Nowadays, there are mainly three basic formulations of diffusion models: denoising diffusion probabilistic models (DDPMs) [88, 93], score matching diffusion models [94, 95], and score SDEs [96, 97].
For DDPMs, the forward and reverse processes are two Markov chains: a forward chain that adds random noise to data and a reverse chain that transforms noise back into data. Formally, denote the data distribution as $\boldsymbol{x}_{0}\sim q(\boldsymbol{x}_{0})$, the forward Markov process gradually adds Gaussian noise to the data according to transition kernel $q(\boldsymbol{x}_{t}|\boldsymbol{x}_{t-1})$. It generates a sequence of random variables $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots,\boldsymbol{x}_{T}$. Thus the joint distribution of $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots,\boldsymbol{x}_{T}$ conditioned on $\boldsymbol{x}_{0}$ is
@@
where $\theta$ is the model parameters and $p_{\theta}(\boldsymbol{x_{t-1}|x_{t}})=\mathcal{N}(\boldsymbol{x}_{t-1};\mu_{\theta}(\boldsymbol{x}_{t},t),\sum_{\theta}(\boldsymbol{x}_{t},t)).$ The key to achieving the success of sample generating is training the parameters $\theta$ to match the actual reverse process, that is, minimizing the Kullback-Leibler divergence between the two joint distributions. Thus, according to Jensen’s inequality, the training loss is
@@
$\geq\mathbf{E}[-\log p_{\theta}(\boldsymbol{x}_{0})]+const.$
For score-based diffusion models, the key idea is to perturb data with a sequence of Gaussian noise and then jointly estimate the score functions for all noisy data distributions by training a deep neural network conditioned on noise levels. The motivation of the idea is that, in many situations, it is easier to model and estimate the score function than the original probability density function. Langevin dynamics is one of the proper techniques. With a step size $\alpha>0$, the number of iterations $T$, and an initial sample $x_{0}$, Langevin dynamics iteratively does the following estimation to gain a close approximation of $p(\boldsymbol{x})$
@@
For score SDEs, the diffusion operation is processed according to the stochastic differential equation (SDE) [97]:
@@
where $\delta(\frac{t}{T})=\delta_{t}$ as T goes to infinity. With any diffusion process in the form of (16), the reverse process can be gained by solving the following SDE:
@@
4 Contrastive-based Methods
Contrastive learning is a widely used self-supervised learning strategy, showing a strong learning ability in computer vision and natural language processing. Unlike discriminative models that learn a mapping rule to true labels and generative models that try to reconstruct inputs, contrastive-based methods aim to learn data representations by contrasting between positive and negative samples. Specifically, positive samples should have similar representations, while negative samples have different representations. Therefore, the selection of positive samples and negative samples is very important to contrastive-based methods. This section sorts out and summarizes the existing contrastive-based methods in time series modeling according to the selection of positive and negative samples. The illustration of the contrastive-based SSL for time series is shown in Fig. 3. In Appendix E.1 - E.5, the main advantages and disadvantages of five contrastive-based submethods are summarized.
@@ 4.1 Sampling contrast
where $S(\cdot)$ denotes the similarity of the two representations. However, due to the non-stationary characteristics of most time series data, it is still a challenge to choose the correct positive and negative samples based on contextual information in time series data. Temporal neighborhood coding (TNC) was recently proposed to deal with this problem [106]. TNC uses augmented Dickey-Fuller (ADF) statistical test to determine the stationary region and introduces positive-unlabeled (PU) learning to handle the problem of sampling bias by treating negative samples as unknown samples and then assigning weights to these samples. The learning objective is extended to
@@
where $w$ is the probability of sampling false negative samples, $\mathcal{N}$ denotes the neighboring area, and $\tilde{\mathcal{N}}$ denotes the non-neighboring area. Supervised contrastive learning (SCL) [107] effectively addresses sampling bias, so introducing the supervised signal to identify positive and negative samples is a feasible solution. Neighborhood contrastive learning (NCL) is a recent time series modeling method that combines context sampling and the supervised signal to generate positive and negative samples [108]. NCL assumes that if two samples share some predefined attributes, then they are considered to share the same neighboring area.
@@ 4.2 Prediction contrast
In this category, prediction tasks that use the context (present) to predict the target (future information) are considered self-supervised pretext tasks, and the goal is to maximally preserve the mutual information of the context and the target. Contrastive predictive coding (CPC) proposed by [109] provides a contrastive learning framework to perform the prediction task using InfoNCE loss. As shown in Fig. 3(b), the context $c_{t}$ and the sample from $p(x_{t+k}|c_{t})$ constructs positive pairs, and the samples from the ‘proposal’ distribution $p(x_{t+k})$ are negative samples. The learning objective is as follows:
@@ 4.3 Augmentation contrast
Since time series data can be converted to frequency domain representations through Fourier transform, the augmentation method can be developed from the time and frequency domains. In the time domain, TS-TCC[116] and its extended version CA-TCC[117] designed two time series data augmentation techniques, one is strong augmentation (permutation-and-jitter), and the other is weak augmentation (jitter-and-scale). TS2Vec [118] generates different views through masking operations that randomly mask out some time steps. Generally speaking, there is no one-size-fits-all answer to the choice of data augmentation methods. Therefore, some works comprehensively compare and study the augmentation methods and further evaluate the performance on different tasks [119, 45, 120, 121]. All the above methods only need a single time series sample in the augmentation operation, while Mixing-up [122] fuses two time series samples to generate a newly augmented view, while the pretext task is to correctly predict the proportion of two original time series samples in augmented view.
Data augmentation in the frequency domain is also feasible for time series data. CoST [123] is a disentangled seasonal-trend representation learning method, which uses fast Fourier transform to convert different augmented views into amplitude and phase representations, and then uses (25) to train the model. BTSF [124] is a contrastive-based method based on a time-frequency fusion strategy, which first generates an augmented view in the time domain through the dropout operation and then generates another augmented view in the frequency domain through Fourier transform. Finally, the bilinear temporal-spectral fusion mechanism is used to achieve the fusion of time-frequency information. However, CoST and BTSF do not modify the frequency representation, while TF-C [125] directly augments the time series data through frequency perturbations, which has achieved better performance than TS2Vec [118] and TS-TCC [116]. Specifically, TF-C implements three augmentation strategies: low- vs. high-band perturbations, single- vs. multi-component perturbations, and random vs. distributional perturbations.
@@ 4.4 Prototype contrast
In time series modeling based on prototypes contrast, ShapeNet [136] takes shapelets as input and constructs a cluster-level triplet loss, which considers the distance between the anchor and multiple positive (negative) samples as well as the distance between positive (negative) samples. ShapeNet is an implicit prototype contrast because it does not introduce explicit prototypes (cluster centers) during the training phase. TapNet [137] and DVSL [138] are explicit prototypes contrast because explicit prototypes are introduced. TapNet introduces a learnable prototype for each predefined class and classifies the input time series sample according to the distance between the sample and each class prototype. DVSL defines virtual sequences, which have the same function as prototypes, i.e., minimize the distance between samples and virtual sequences, but maximize the distance between virtual sequences. MHCCL [139] proposes a hierarchical clustering based on the upward masking strategy and a contrastive pairs selection strategy based on the downward masking strategy. In the upward mask strategy, MHCCL believes that outliers greatly impact prototypes, so these outliers should be removed when updating prototypes. The downward masking strategy, in turn, uses the clustering results to select positive and negative samples, i.e., samples belonging to the same prototype are regarded as true positive samples, and samples belonging to different prototypes are regarded as true negative samples.
@@ 4.5 Expert knowledge contrast
Expert knowledge contrast is a relatively new representation learning framework. Generally speaking, this modeling framework incorporates expert prior knowledge or information into deep neural networks to guide model training [140, 141]. In the contrastive learning framework, prior knowledge can help the model choose the correct positive and negative samples during training. An example of expert knowledge contrast is shown in Fig. 3(e).
Here we sort out three typical works of expert knowledge contrast for time series data. Shi et al. [142] used the DTW distance of time series samples as prior information and believed that two samples with a small DTW distance have a higher similarity. Specifically, given the anchor $x^{ref}$ and the other two samples $x_{i}$ and $x_{j}$, the DTW distance between $x^{ref}$ and the other two samples is calculated first, then the sample with a small distance from $x^{ref}$ is considered as the positive sample of $x^{ref}$. This selection process is defined as
@@
Actually, introducing more prior knowledge in contrastive-based SSL can help the model extract better representations. The trend of this family of methods can be summarized from two perspectives: (i) Addressing sampling bias. Sampling bias is caused by inappropriate selection of positive and negative samples, so introducing prior knowledge useful for selecting positive and negative samples can deal with this problem, such as a clustering-based negative sample detection algorithm [145] and sample identification strategy based on real labels [107, 108]. (ii) Addressing representation bias. Representation bias means that the extracted representations cannot be guaranteed to be strongly related to the downstream task. The essential reason is that there may be a big difference between the goals of the pretext task and the downstream task. An interesting trend is to fuse semi-supervised learning and contrastive-based SSL to guide the training of the encoder through a small amount of labeled data [146, 147].
@@ 5 Adversarial-based Methods
According to the final task, the existing adversarial-based representation learning methods can be divided into time series generation and imputation, and auxiliary representation enhancement. The illustration of the adversarial-based SSL for time series is shown in Fig. 4. In Appendix F.1 - F.2, the main advantages and disadvantages of two adversarial-based submethods are summarized. Furthermore, the main differences in characteristics and limitations between the adversarial-based methods and the previous two methods (generative-based and contrastive-based) are shown in Appendix G.
@@ 6 Applications and Datasets
TABLE II: Summary of time series applications and widely used datasets. The UCR and UEA datasets have multiple sub-datasets respectively, so their size and dimension are not fixed values. The size and dimension of each sub-dataset are represented by $\mathcal{M}$ and $\mathcal{D}$, respectively. AnRa represents Anomaly Ratio. SaIn represents Sampling Interval. Datasets for classification tasks and clustering tasks are listed together because the goals of these two tasks are similar. AD represents anomaly detection. F represents forecasting. C&\&&C represents classification and clustering.
App Dataset Size Dim Comment
@@
SSL has many applications across different time series tasks. This section summarizes the most widely used datasets and representative references according to the application area, including anomaly detection, forecasting, classification, and clustering. As shown in Table II, we provide useful information, including dataset name, dimension, size, source, and useful comments. For each task, we summarize from the following aspects: task description, related methods, evaluation metrics, examples, and task flow. Due to space limitations, relevant descriptions of evaluation metrics, examples, and task flow can be found in Appendix H. In addition, we provide performance comparison results of different methods on the same dataset and further summarize the correlation between methods and tasks, the details can also be found in Appendix I.
6.1 Anomaly detection
  • Task description. The anomaly detection problem for time series is usually formulated as identifying outlier time points or unexpected time sequences relative to some norm or usual signal.
@@ 6.1 Anomaly detection
  • Related methods. Most time series anomaly detection methods are constructed under an unsupervised learning framework because obtaining labels for anomalous data is challenging. Autoregressive-based forecasting and autoencoder-based reconstruction are the most commonly used modeling strategies. To be concrete, THOC [52] and GDN [55] employ autoregressive-based forecasting SSL framework, which assumes that anomalous sequences or time points are not predictable. RANSynCoders [63], USAD [66], AnomalyTrans [163], and DAEMON [184] employ autoencoder-based reconstruction SSL framework. Furthermore, VGCRN [83] and FuSAGNet [67] combine two frameworks to achieve more robust and accurate results. It is beneficial to introduce an adversarial-based SSL, which can further amplify the difference between normal and anomalous data, such as USAD [66] and DAEMON [184].
6.2 Forecasting
  • Task description. Time series forecasting is the process of analyzing time series data using statistics and modeling to make predictions of future windows or time points.